A sample of dolomite contains 45% of chcl3 40% of MGSU 3 and 15% play calculate mass of H2 S o4 of 30% Jain required to react completely with 10 gram sample

Grade:11

1 Answers

Arun
5 years ago
Supposing clay doesn't react with H2SO4. 
Supposing the given percentages are by mass: 

CaCO3 + H2SO4 → CaSO4 + H2O + CO2 
(45% of 10 g) / (100.0875 g CaCO3/mol) x (1 mol H2SO4 / 1 mol CaCO3) = 
0.04496 mol H2SO4 to react with the CaCO3 

MgCO3 + H2SO4 → MgSO4 + H2O + CO2 
(40% of 10 g) / (84.31413 g MgCO3/mol) x (1 mol H2SO4 / 1 mol MgCO3) = 
0.04744 mol H2SO4 to react with the MgCO3 

(0.04496 mol + 0.04744 mol) x (98.0791 g H2SO4/mol) / (0.30) = 30 g 30% H2SO4